OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y?

OpenStudy (anonymous):

Answer choices available: A. 3 B. 0 C. 6 D. -2

OpenStudy (anonymous):

From the first equation, we can rearrange to get x: x - y = 3 x = y + 3 Plug this into the second equation and we get: 3(y + 3) + 8y = 9 3y + 9 + 8y = 9 11y = 0 y = 0 Then substitute y back in the original equation: x - y = 3 x - 0 = 3 x = 3 A is your answer.

OpenStudy (anonymous):

Thank you SirSushii! :)

OpenStudy (mathmate):

What is the value of y? 3 times the first equation subtract the second gives 3x - 3y = 9 and 3x - 8y = 9 Subtracting, 5y=0 => y=0 B is the answer for y.

OpenStudy (anonymous):

Oops. My bad. xD

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

Oh dang! Well, it is the thought that counts! :) Thanks mathmate, and I still thank you SirSushii for trying to help me! Thanks both of you! :)

OpenStudy (mathmate):

You're welcome! :)